# Solving Sequential Greedy Problems Distributedly with Sub-Logarithmic Energy Cost

### v1

@@ Abstract
We study the awake complexity of graph problems that belong to the class O-LOCAL, which includes a large subset of problems solvable by sequential greedy algorithms, such as $(\Delta+1)$-coloring, maximal independent set, maximal matching, etc. It is known from previous work that, in $n$-node graphs of maximum degree $\Delta$, any problem in the class O-LOCAL can be solved by a deterministic distributed algorithm with awake complexity $O(\log\Delta+\log^{\star}n)$.
@@ 1 Introduction
This paper focuses on the awake complexity of solving a large class of natural problems, namely those in the class O-LOCAL, introduced in [BM21]. Roughly, this class includes all problems which can be solved by a restricted form of sequential greedy algorithms. More specifically, the order in which the nodes are sequentially picked is governed by an arbitrary given acyclic orientation of the edges of the graph $G$, that is, a node $v$ can be picked only if solutions have been computed for all its descendants according to the given orientation. More importantly, the solution at $v$ can be computed based only on the solutions previously computed for these descendants. Even if O-LOCAL does not contain all sequentially greedily solvable graph problems (e.g., distance-2 coloring is not in O-LOCAL), it includes many among the most important problems in distributed network computing, such as, $(\Delta+1)$-vertex coloring, maximal independent set (MIS), maximal matching, $(2\Delta-1)$-edge coloring, etc, where $\Delta$ denotes the maximum degree in the graph.
@@ 1.2 Additional Related Work
The Sleeping model was introduced in [CGP20], which opened a large avenue of research dedicated to measuring the potential benefit of providing the nodes of a distributed system with the ability to turn on and off at will. In particular, MIS has attracted a lot of attention in this context (see [CGP20, GP22, GP23, HPR22]), and it was eventually shown [DJP23] that MIS can be solved (by a randomized algorithm) with awake complexity $O(\log\log n)$. Regarding $(\Delta+1)$-coloring, the aforementioned paper [BM21] on solving O-LOCAL problems established a (deterministic) awake complexity $O(\log\Delta+\log^{\star}n)$. The randomized algorithm in [DFRZ24] has awake complexity $O(\log\log n)$. This presents an improvement over the best known randomized algorithms for $(\Delta+1)$-coloring in the LOCAL model [CLP18, HKNT22], which have round complexity $\tilde{O}(\log^{2}\log n)$, though to the expense of a significant increase of the round complexity. Several other problems have been studied in the Sleeping model, including matching [GP22], vertex cover [GP22], spanning tree [BM21], and minimum-weight spanning tree [AJP24]. Note that problems such as maximal matching and $(2\Delta-1)$-edge-coloring are in O-LOCAL, and can thus be solved with awake complexity $O(\log\Delta+\log^{\star}n)$ [BM21]. A direct consequence of the result in this paper is that MIS, $(\Delta+1)$-coloring, maximal matching, and $(2\Delta-1)$-edge-coloring have awake complexity $O(\sqrt{\log n}\cdot\log^{\star}n)$.
@@ 6 Conclusion and Open Problems
Finally, understanding how much randomization can help is also an interesting open question.

### v2

@@ Abstract
We study the awake complexity of graph problems that belong to the class O-LOCAL, which includes a subset of problems solvable by sequential greedy algorithms, such as $(\Delta+1)$-coloring and maximal independent set. It is known from previous work that, in $n$-node graphs of maximum degree $\Delta$, any problem in the class O-LOCAL can be solved by a deterministic distributed algorithm with awake complexity $O(\log\Delta+\log^{\star}n)$.
@@ 1 Introduction
This paper focuses on the awake complexity of solving a large class of natural problems, namely those in the class O-LOCAL, introduced in [BM21]. Roughly, this class includes all problems which can be solved by a restricted form of sequential greedy algorithms. More specifically, the order in which the nodes are sequentially picked is governed by an arbitrary given acyclic orientation of the edges of the graph $G$, that is, a node $v$ can be picked only if solutions have been computed for all its descendants according to the given orientation. More importantly, the solution at $v$ can be computed based only on the solutions previously computed for these descendants. Even if O-LOCAL does not contain all sequentially greedily solvable graph problems (e.g., distance-2 coloring is not in O-LOCAL), it includes important problems in distributed network computing, such as $(\Delta+1)$-vertex coloring and maximal independent set (MIS), where $\Delta$ denotes the maximum degree in the graph.
@@ 1.2 Additional Related Work
The Sleeping model was introduced in [CGP20], which opened a large avenue of research dedicated to measuring the potential benefit of providing the nodes of a distributed system with the ability to turn on and off at will. In particular, MIS and $(\Delta+1)$-coloring have attracted a lot of attention in this context (see [CGP20, GP22, GP23, HPR22, DFRZ24, BM21]). The aforementioned paper [BM21] on solving O-LOCAL problems established that both $(\Delta+1)$-coloring and MIS have deterministic awake complexity $O(\log\Delta+\log^{\star}n)$. A direct consequence of the result in our paper is that MIS and $(\Delta+1)$-coloring have deterministic awake complexity $O(\sqrt{\log n}\cdot\log^{\star}n)$. Regarding randomized algorithms, it was shown that MIS and $(\Delta+1)$-coloring have awake complexity $O(\log\log n)$ [DJP23, DFRZ24]. This presents an improvement over the best known randomized algorithms for $(\Delta+1)$-coloring in the LOCAL model [CLP18, HKNT22], which have round complexity $\tilde{O}(\log^{2}\log n)$, though to the expense of a significant increase of the round complexity. Several other problems have been studied in the Sleeping model, including matching [GP22], vertex cover [GP22], spanning tree [BM21], and minimum-weight spanning tree [AJP24].
@@ 6 Conclusion and Open Problems
Moreover, understanding how much randomization can help is also an interesting open question.
@@
Finally, understanding whether our techniques can be applied to solve edge problems such as maximal matching and $(2\Delta-1)$-edge coloring remains an interesting open question.
Open Question 5.
Can we extend O-LOCAL to include edge problems and solve them with deterministic sublogarithmic awake complexity?
Acknowledgements
We thank William K. Moses Jr. for pointing out that maximal matching does not belong to the class O-LOCAL.